AO Commissioning on 23 June 2011

Observer(s):S. Esposito, JC. Guerra, F. Quiros-Pacheo, J. Argomedo, D.Miller, D. Thompson

Telescope Operator: Geno (LBTO)

SW Support: N. Cushing (Tucson)

Telescope Support: J. Urban, J. Little (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Details

2:00 Trouble with a breaker on the right chamber door. John and Jeff fix is and get the door open

2:15 Juan Carlos start taking sky flats Br-g, H2, FeII, Ks H, J

2:48 iif_preset, M5_1395, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • Failed, M2 Hexapod not home. The AO Guys had turned off power, and told us, so we need to re-home
  • Hexapod rehomed

2:54 iif_preset, M5_1395, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • Success, but no star
  • Geno clear C00, Adaptive Optics on Primary
  • No star
  • Spiral search
  • Star found and centered

3:20 iif_preset, SDT_Mkn231, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - Success
  • First SVT target
  • On-axis star found

3:23 iif_preset, SDT_Mkn231, MODE=ACTIVE, GS=0, PA=90
  • Collimate on-axis first
  • Guide star has axis ratio of 6to1
  • Iterate 8 times to get well collimated telescope

3:30 observe, Mkn231_test
  • Failed - had PA=90 and could not reach Guide Star

3:31 observe, Mkn231_test
  • Set PA=100, Could reach it

The TT offlading are been refuced. The secondary shows about 10 umPV

4:20 Open loop and go to Zenith to turn off URTH fans

4:30 Restart PCS because we change PCSInstrument.conf back to RightPointingOrigin to 0.0
  • had set it to 0.6, -4.8 to try to put on_axis targets on the middle of the Pisces
  • in earlier presets, on-azis was in the same position as before, with 0,0 so this parameter seems to not work

4:45 GCS got stuck with the error "resetting exposure time". Tried to kill GCS, restart, next preset stuck, authorize stuck, restart IIF, kill GCS, authorize complete....

5:00 Preset in ACQUIRE mode, Geno set IE, CA

5:05 ACE Preset to Mkn231 but no guide star seen on Acquisition image

5:15 Back on-axis OK

5:10 ACE Preset, Reference star is on AO TV, but nothing on GCS

5:15 Same preset as above, but with PA=110 in stead of 100. No Guide star on AO TV or Guider

5:30 Acquire Preset to QSO2198_16. No star on guider, but target is 16 mag. Spiral search see a dim star.

5:40 Acquire Preset to TS=1, but nothing. Spiral search, but nothing.

5:45 Acquire Preset to ACT0328. Clear all Active Optics and C00.
  • spiral search and finally find a star
  • Geno set IE, CA

6:00 Acquire preset to BS9168 to collimate. This is a trusted star.

6:08 Active Preset. Collimate

6:15 Acquire Preset to QSO2198. Target and reference stars seen on guider. Geno center

6:21 Observe, SDT_QSO_2198_16
# Test of observing script
# 27Apr2011  DLM
preset, OBJECT=SDT_QSO2198_16, MODE=ace, GS=5, RS=0, TS=1, PA=90
auto_collimate
runao
set_position
camera, EXP=10.0, FRAMES=1, COADDS=1, FILTER=5
getimage, COMMENT=Test
camera, EXP=1.0, FRAMES=1, COADDS=1, FILTER=5
getimage, COMMENT=Test

6:45 observe, file3, file2, /ACQUIRE
-----------------------------------------------------------------------------------------
   => Format correct in script /OldRepository/PISCES_Data/20110623/scripts/QSO2198.obs
   => Flag values acceptable in script /OldRepository/PISCES_Data/20110623/scripts/QSO2198.obs
----------------------------------------------------------------------------------------------
  
      # Target star: 0   SDT_QSO2198_16
        16:02:10.1920 +30:47:55.520 RADEC 2000.0 J2000 0.0 0.0 16.00 R 0.00 B_V 1.65
      # Guide star: 0
        16:02:10.1920 +30:47:55.520 RADEC 2000.0 J2000 0.0 0.0 16.00 R 0.00 B_V 0.63
  
      preset, SDT_QSO2198_16, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  
*** Waiting for the user to press the "OK" button on the popup box ***
  • Geno positions the on-axis star.
  • Out of focus guiding star.

6:57 Ace preset failed because Guider choose ghost of highly out of focus guide stat.

6:58 Acquire preset to set better focus by hand

7:01 Observe the QSO2198_16 with a dither pattern.
  • Good.

7:52 Observe the PSF star. preset, PSF_QSO2198, MODE=ACE, GS=5, RS=0, PA=30.0000
  • Cannot see the PSF star. Change target.

8:10 Observing planet transit at 89 degrees elevation. Sent preset and close the loop. The observing technique requires to stop the rotator.
  • About 40 minutes close loop the shell rips " too much stroke in the secondary". Possible due to Azimuth stopped.

8:51 Sent the preset again. .....Failed. the azimuth stopped. Gino recover it.
  • Sent again the Observe script. And Simone and Gino went to check the three house temperature and building/azimuth
  • Stop observing. Telescope locked

8:57. Simone went to check the three house and Gino goes to check building and azimuth.
  • Simone informed that the temperature in the three house is about 23 degrees.

9:20 Restart the telescope..

9:21 iif_preset, SDT_BD303639, MODE=ACQUIRE,GS=0,ROTMOMODE=IDLE.
  • observe.....
  • sent the preset, ref acquire and was starting to acquire the reference.
  • Azimuth stopped. HBS problem.

9:34 Sent again.
  • Observe....
  • close loop... Automatic gain procedure
  • taking dithering pattern.

10:05 iif_preset,. SDT_HB8779, MODE=ACQUIRE,GS=0,ROTMODE=IDLE.
  • point the telescope, align to the GS and to check the rotation position.

  • Observe script.
  • preset, SDT_HB8779 ,MODE=ACE,GS=6,RS=0,PA=-90
    • collimation..
    • acquire ref.
    • close loop failed. Close the loop from the AOS.
    • Problem with a high noise in the CCD39 at bin 1, this stop the loop. Manually changing from bin 1 to 2 solve the problem. The AO system can be recovered success and at this point the HBS failed. Open the loop. Stop the observations.

10:37 Resend the preset. iif_preset, SDT_HB8779, MODE=ACQUIRE,GS=0, ROTMODE=IDLE.
  • Failed. GS problem, restart the GCS system.

10:43 Resend the preset. iif_preset, SDT_HB8779, MODE=ACQUIRE,GS=0, ROTMODE=IDLE.
  • Failed. HBS failed.

10:47 Resend the preset. iif_preset, SDT_HB8779, MODE=ACQUIRE,GS=0, ROTMODE=IDLE.
  • Failed. HBS failed.


STOP THE OBSERVATIONS............ HBS NOT RECOVERED.

Summary of darks and flats

darks: pisces.20110622.0002->0076
flats: 0 (2.14) 0085->0101
  Br_gm 0102->0122
  H2 0123->0128, 0131->0139
  FeII 0144->0154, 209->0219
  Ks 0156->0175
  H 0177->0186, 0221->0230, 0243->0252
  H2 0187->0197
  J 0198->208, 0232->0242
  FeII

-- DougMiller - 23 Jun 2011
Topic revision: r9 - 10 Feb 2012, MichaelWagner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback